OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x − 3y = 3 5x − 4y = 4 −1 0 x can be any number as there are infinitely many solutions to this system There is no x value as there is no solution to this system

OpenStudy (anonymous):

2x − 3y = 3 * 5 ----------- i 5x − 4y = 4 * -2 ----------- ii 10x - 15y = 15 -10x + 8y = -8 adding the two - 7y = 7 y = -1 substitute in i 2x - 3(-1) = 3 2x + 3 = 3 2x = 3-3 2x = 0 x = 0 so solution (0, -1) so x = 0